40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1445     Entry time: Mon Mar 30 15:51:27 2009
Author: steve 
Type: Update 
Category: Electronics 
Subject: HP4291A left the lab to be repaired 

Eric Gustafson is handling the old HP4291A rehabilitation. Tarac picked both units up today.

March of 2008 Tucker Electronics failed to fix it's intermittent ~25MHz 0.5V oscillation at the swept sine output

See 40m-elog id:398 on 3-24-2008 by Rob Ward

 

ELOG V3.1.3-